OpenStudy (anonymous):

solve the equation in the specified interval. 1. sec x=-3, -π≤x<π 2. sin x=-0.5, -∞

OpenStudy (anonymous):

secx=-3 1/cosx = -3 -1/3 = cosx now -1/3 is about cos(109.47) so: cosx = cos(109.47) so x1 = 109.47 + 360k and x2 = -109.47 + 360k x1: we can take only k=0 so 109.47 and x2 we can take only k=0 so -109.47 two solutions : 109.47,-109.47